3. What is the future value of $4000 in 5 years if it is invested at an interest rate of 6.00%? a. $5,300 b. $5,563 c. $5,353 d. $6,214 e. $6,431
Answer #1

Assuming that the Money is compounded annually- this is important since compounding frequency will have an impact on the future value.

The formula to calculate future value is :

FV = P ( 1 +i)^n

In our case ,

P = 4000

i = 6%

and n = 5 yrs

Using the values in above equation, FV = 4000 (1+ 6% )^5 = 5,352.90 = 5,353